OpenStudy (anonymous):

3:12 = x:16 solve for x

OpenStudy (australopithecus):

3/12 = x/16 3(16)/12 = x

OpenStudy (anonymous):

 You cross multiply 3 and 16=48, then divide by 12 which then equals 4. :) x=4